OpenStudy (anonymous):

The surface area of a spherical balloon is 175 cm2. If the length of the radius was six times as large, what would the surface area be?

OpenStudy (cornzyblack):

Since the Surface area was 175 cm^2 , we use the foemula to calculate for the radius(making r subject of formula) Surface area of sphere = 4 * pi * r^2 So r = Sqrt(Surface area of sphere / pi ) So r = 7.463 if the radius was 6 times as large then, new radius R = r * 6 = 44.778 Using Surface area formula = 4 * pi * R ^2 New Surface area is , 25,196.44 cm^2

ganeshie8 (ganeshie8):

@CornzyBlack method works always.. just he forgot about "4".. you should get a surface area : 175 x 6^2 = 6300 sq.cm

OpenStudy (anonymous):

Well the answer choices are: 1050cm^2 6300cm^2 8400cm^2 11200cm^2

OpenStudy (anonymous):

Oh, ok thanks you guys!

ganeshie8 (ganeshie8):

yw..